In March 2004, vegas.com gave odds on who would win the 2004 World Series. They gave the New York Yankees 2 to 1 odds and gave the Boston Red Sox 5 to 2 odds. (these are odds against)
a.) convert both odds to the probability of winning.
b.) who did they think was more like to win and why?
c.) If you had bet $10 on the Boston Red Sox, and they went on to win the World Series, how much money did you win, in addition to having your wager returned?

Answer by stanbon(75887)

In March 2004, vegas.com gave odds on who would win the 2004 World Series. They gave the New York Yankees 2 to 1 odds and gave the Boston Red Sox 5 to 2 odds. (these are odds against)
---
Fact: odds against E = [P(not E)]/P(E)]
-----------------------------------------------
a.) convert both odds to the probability of winning.
NY Yanks:
odds against = 2:1
P(not winning) = 2/3
P(winning) = 1/3
----
Boston:
odds against = 5:2
P(not win) = 5/7
P(win) = 2/7
--------
---------------------------
b.) who did they think was more like to win and why?
Compare 1/3 and 2/7
Cross-multiply to get: 7 ? 6
---
7>6, so 1/3 > 2/7
They favor the Red Sox.
============================
c.) If you had bet $10 on the Boston Red Sox, and they went on to win the World Series, how much money did you win, in addition to having your wager returned?
---
I'm not familiar with how odds pay off.
==========================================
Cheers,
Stan H.